OpenStudy (anonymous):

Evaluate. 1 – 3 • 7 + 4 • 2 A. –20 B. –28 C. –12 D. –32

OpenStudy (anonymous):

It's C -12

OpenStudy (welshfella):

@angelierivera - please explain how you got that

OpenStudy (anonymous):

If you want help in math you can get an app on your phone called Wolfram Alpha it helps with all subjects with the answers right there but it's just for a small price

OpenStudy (anonymous):

I got it by asking Siri

OpenStudy (anonymous):

@welshfella

OpenStudy (ddos_dragon):

1 – 3 • 7 + 4 • 2 1- 21 + 4 * 2 1 - 21 + 8 - 20 + 8 -12

OpenStudy (koalagurl1):

She's right. It is C.

OpenStudy (welshfella):

please read the code of conduct - we are here to show people how to get answers not just give them

OpenStudy (ddos_dragon):

Remember to medal and close the question. Have a nice day!

OpenStudy (ddos_dragon):

Indeed, they will just ask again. The purpose of this site is to further the understanding of those who require assistance.

OpenStudy (welshfella):

right

OpenStudy (koalagurl1):

You multiply first so its 3*7 and 4*2

OpenStudy (koalagurl1):

21 and 8

OpenStudy (ddos_dragon):

By the way, you must familiarize yourself with PEMDAS

OpenStudy (anonymous):

Sorry but if you want to learn the steps get the app Wolfram alpha

OpenStudy (ddos_dragon):

Parentheses Exponent Multiplication & Division Addition & Subtraction

OpenStudy (anonymous):

See it will give you step by step solutions!

OpenStudy (ddos_dragon):

When solving equations with multiple operations, do the operations in that order ^

OpenStudy (koalagurl1):

so its 1-21= -20

OpenStudy (ddos_dragon):

-20 + 8 = -12 :)

OpenStudy (ddos_dragon):

Remember to close your question.

OpenStudy (koalagurl1):

and -20 + 8 = -12

OpenStudy (ddos_dragon):

@joyyy12 Close the question :P

OpenStudy (koalagurl1):

See, @angelierivera is right. It's C.

OpenStudy (ddos_dragon):

She never doubted the answer, she just stated that you have to walk the asker through.

OpenStudy (koalagurl1):

I did.

OpenStudy (anonymous):

Hey can any of you guys help me on my question??

OpenStudy (ddos_dragon):

You cannot just blatanly give out the answers, you have to explain it to whoever asked :P

OpenStudy (ddos_dragon):

Tag me.

OpenStudy (ddos_dragon):

*Blatantly

OpenStudy (anonymous):

My orevious question?

OpenStudy (koalagurl1):

I DID.

OpenStudy (ddos_dragon):

Tag me :P

OpenStudy (koalagurl1):

Have you not read the comments?? @Ddos_Dragon

OpenStudy (koalagurl1):

I explained it.

OpenStudy (ddos_dragon):

I read the comments :P

OpenStudy (koalagurl1):

Ugh. Whatever. :|

OpenStudy (ddos_dragon):

I was just explaining to you, that welsh was not doubting angel's answer she was just informing angel that he has to explain how he got there to the asker!

OpenStudy (koalagurl1):

idc.

OpenStudy (koalagurl1):

I explained it though.

OpenStudy (koalagurl1):

Also, yes. :) @angelierivera

OpenStudy (koalagurl1):

Just tag me. @angelierivera